Identify the atom with the following ground-state electron configuration for its valence shell.

4s2 3d10
Answer #1
Concepts and reason

Electron configuration of an atomic species describes the properties of atoms and helps us to understand the basic properties of that atom; like shape of an atom, energy of its electron and also other properties like bonding behavior of atoms.

For the electron configuration there are many rules or principle to filed out electron in there orbital or sub orbital and electron assigning by the prospective energy state, however these assignment are arbitrary but its help us to better understanding of its bonding ability, magnetism properties and other chemical properties of atoms.

Fundamentals

For the electron configuration of an atomic species, there are mainly three principles or rules:

First is Aufbau’s principle for filled out electron in their respective orbital, filled according to their orbital energy and its also called the building block principle.

Second is Hund’s rule that describe first of all every orbital in a sublevel will be singly field than any orbital will get doubly filed.

Third is Pauli-exclusion principle that describe two identical half-integer spin can’t be filed in same quantum state.

subshell
number of orbitals
total number of possible
electrons in each orbital
1
3(Px,Py, Pz)
5(dx2 - y2, dz2, dxy, dxz, dyz)

By using Aufbau’s principle for electronic configuration electron filed in increasing order of orbital:

1s<2s<2p<3s<3p<4s<3d<4p<5s<4d<5p<6s<4f<5d<6p<7s<5f<6d<7p1s < 2s < 2p < 3s < 3p < 4s < 3d < 4p < 5s < 4d < 5p < 6s < 4f < 5d < 6p < 7s < 5f < 6d < 7p For the electronic configuration of 4s23d104{s^2}3{d^{10}}

So, start the filling one by one:

4s23d10=1s2<2s2<2p6<3s2<3p6<4s2<3d104{s^2}3{d^{10}} = 1{s^2} < 2{s^2} < 2{p^6} < 3{s^2} < 3{p^6} < 4{s^2} < 3{d^{10}}

Total number of electron in this configuration = 2 + 2 + 6 + 2 + 6 + 2 + 10 = 30

So, the total electron in ground state is 30 and the atomic number also will be 30 so,

30 is the atomic number of Zinc

Ans:

Zinc = Zn = electronic configuration in ground state = 4s23d104{s^2}3{d^{10}}